Lakeside Inc. produces Product A and Product B that require special machining time. Machine time capacity is 16,970 machine hours per month. Lakeside estimates April demand for each product and provides additional information as follows:

Product A Product B April demand 4,700 units 3,700 units Contribution margin $ 753 per unit $ 670 per unit Required machine hours 3 per unit 2 per unit Required:

Determine how many units of each product Lakeside should produce to maximize contribution margin in April.
Answer #1

Though the contribution margin for product A is more than product B but the limiting factor is machine hours product A uses more machine hours as compared to to product B.

Therefore we have to calculate contribution per machine hour which will be

Product A = 16970/(753*3)= 7.51

Product B = 16970/(670*2)= 12.66

So so as the contribution per machine hour is more of product B lakeside should produce maximum of product B and then the remaining machine hours should be allocated to product A

Units produced

Product B = 3700 units

Product A

Machine hours left = 16970 - (3700*2) = 9570

Products = 9570/3 = 3190 units

Total contribution = 3700*670 + 3190*753

=$ 4,881,070

If you find the answer helpful please upvote